# Histopathology Foundation Models Enable Accurate Ovarian Cancer Subtype Classification

### v1

@@ Abstract
Large pretrained transformers are increasingly being developed as generalised ‘foundation models’ which can underpin powerful task-specific artificial intelligence models. Histopathology foundation models show great promise across many tasks, but analyses have typically been limited by arbitrary hyperparameters that were not tuned to the specific task/dataset. We report the most rigorous single-task validation conducted to date of a histopathology foundation model, and the first performed in the context of ovarian cancer morphological subtyping. Attention-based multiple instance learning classifiers were compared using vision transformer and ResNet features generated through varied preprocessing and pretraining procedures. The training set consisted of 1864 whole slide images from 434 ovarian carcinoma cases at Leeds Teaching Hospitals NHS Trust. Five-class classification performance was evaluated using the F1 score, AUROC, and balanced accuracy through five-fold cross-validation, and these cross-validation models were ensembled for evaluation on a hold-out test set and an external set from the Transcanadian study. Reporting followed the TRIPOD+AI checklist. The vision transformer-based histopathology foundation model, UNI, performed best in every evaluation, with five-class balanced accuracies of 88% (95% CI: 82-94%) and 93% (87-98%) in hold-out internal and external testing, compared to the best ResNet model scores of 68% (60-76%) and 81% (70-90%), respectively. Normalisations and augmentations aided the generalisability of ResNet-based models, but these still did not match the performance of UNI, which gave the best external performance in any ovarian cancer subtyping study to date. UNI also gave significant improvements over an ImageNet-pretrained transformer in eight out of nine evaluations (p-values 0.0005-0.0306). Histopathology foundation models offer a clear benefit to subtyping, improving classification performance to a degree where clinical utility is tangible, albeit with an increased computational burden. Such models could provide a second opinion to histopathologists diagnosing challenging cases and may improve the accuracy, objectivity, and efficiency of pathological diagnoses overall.
Key Words
@@ Introduction
Histological subtyping is an essential component of the diagnostic process, but it can be challenging. From an individual slide, pathologists only exhibit concordance on an ovarian cancer diagnosis around 80% of the time [6]. As a result, they often request ancillary tests (such as P53 immunohistochemistry) or seek a second opinion from a gynaecological subspeciality expert, with associated logistical and financial burdens. With increasing cancer rates [1] and complexity in diagnostic testing, histopathology services are increasingly struggling to meet demand. For example, most histopathology departments in the UK routinely resort to outsourcing work or hiring temporary staff [7], despite the UK being one of the countries with the most pathologists per capita [8]. Any delays resulting from demand outstripping diagnostic resources risk catastrophic impacts on patient outcomes, with a four-week delay in cancer treatment being associated with an approximately 10% increased mortality rate among patients [9].
Conceptually, artificial intelligence (AI) may offer clinical value by giving an efficient second opinion to histopathologists, streamlining the diagnostic process and perhaps offering additional support when subspecialty experts are not readily available [10]. However, AI models for ovarian cancer diagnosis have not yet demonstrated clinical utility, with most research being small-scale prototyping [11] without regulatory approval for clinical use in Europe or the United States [12]. AI for ovarian cancer subtyping has constituted a small field of research where, aside from our work [13, 14], research has almost exclusively been published by a single group [15, 16, 17, 18, 19, 20, 21]. While the accuracy of such models has increased over time, the best models still only achieve around 80% accuracy [20], and lack sufficient real-life testing.
@@
Recently, some researchers have attempted to create histopathology ‘foundation models’ [30], using self-supervised learning techniques to generate broad histopathological feature sets which are not specific to a single organ/cancer type. These have grown rapidly, from tens of thousands of WSIs used to train models with tens of millions of parameters in 2022 and early 2023 [31, 32, 33, 34, 35, 36], to hundreds of thousands of WSIs and hundreds of millions of parameters more recently [37, 38, 39, 40]. One such model was even trained using 1.5 million WSIs [30]. Foundation models have typically been based on vision transformers (ViTs), utilizing the impressive scalability of transformers seen across many fields, most notably with large language models [41, 42]. Histopathology foundation models have exhibited impressive performance across diverse tasks [34, 39], although analyses have been relatively shallow, without thorough hyperparameter tuning and rigorous statistical comparison of downstream models. Consequently, it is unclear whether models were applied optimally (especially those exhibiting sub-optimal performance), and whether differences between them were significant. Furthermore, many analyses have been conducted using single-centre data, limiting the assessment of models’ generalisability.
@@
In this study, we present the most comprehensive validation conducted to date comparing the standard ResNet50 feature extractor with histopathology foundation models, specifically in the context of ovarian cancer subtyping. This includes comparing whether the performance of ResNet-based MIL classifiers can match those of newer transformer-based MIL classifiers through normalisation, augmentation, and improved tissue detection techniques. The analysis includes rigorous hyperparameter tuning and evaluations through five-fold cross-validation, hold-out testing, and external validation, and was conducted with the largest repository of ovarian cancer WSIs used in any AI study to date.
Methods
Data
A training set of 1864 adnexal tissue WSIs was retrospectively collected from 434 cases of ovarian carcinoma treated at Leeds Teaching Hospitals NHS Trust between 2008 and 2022. Cases were only included if a gynaecological pathologist had diagnosed them as one of the five most common epithelial ovarian cancer subtypes (HGSC, LGSC, CCC, MC, EC). A histopathologist (KA) independently verified all diagnoses, removing any cases with discrepancies. Several representative haematoxylin and eosin (H&E)-stained adnexal tissue glass slides were selected for each case, with only formalin-fixed, paraffin-embedded (FFPE) samples used. These samples were prepared and digitised at 40x magnification using a Leica Aperio AT2 scanner. The population-level class imbalance was reflected in the training set (Table 1), with the least common subtype (LGSC) represented by only 92 WSIs from 21 cases, compared to 1266 WSIs from 308 cases for the most common subtype (HGSC). The training set contained both primary and interval debulking surgery (IDS) specimens, the inclusion of which was previously found to be beneficial to subtype classification [43].
An independent class-balanced hold-out test set was collected through the same protocol, consisting of 100 primary surgery specimen WSIs from 30 patients. An external set of 80 WSIs from 80 patients was accessed from the Transcanadian Study [29]. These had been digitised using an AperioScope scanner and made available at 20x magnification, alongside subtype labels that had been determined by a gynaecological pathologist. This enabled an analysis of generalisability to different slide preparation and scanning procedures.
@@ Methods
Carcinoma Subtype Training WSIs (Patients) Hold-out WSIs (Patients) External WSIs (Patients)
High-Grade Serous (HGSC) 1266 (308) 20 (7) 30 (30)
Low-Grade Serous (LGSC) 92 (21) 20 (6) 9 (9)
Clear Cell (CCC) 198 (45) 20 (7) 20 (20)
Endometrioid (EC) 209 (38) 20 (5) 11 (11)
Mucinous (MC) 99 (22) 20 (5) 10 (10)
Total 1864 (434) 100 (30) 80 (80)

### v2

# A Comprehensive Evaluation of Histopathology Foundation Models for Ovarian Cancer Subtype Classification

@@ Abstract
Large pretrained transformers are increasingly being developed as generalised ‘foundation models’ which can underpin powerful task-specific artificial intelligence models. Histopathology foundation models show great promise across many tasks, but analyses have typically been limited by arbitrary hyperparameters that were not tuned to the specific task/dataset. We report the most rigorous single-task validation of histopathology foundation models to date, specifically in the context of ovarian cancer morphological subtyping. Attention-based multiple instance learning classifiers were compared using three ImageNet-pretrained feature extractors and fourteen histopathology foundation models. The training set consisted of 1864 whole slide images from 434 ovarian carcinoma cases at Leeds Teaching Hospitals NHS Trust. Five-class classification performance was evaluated using balanced accuracy, AUROC, and F1 scores through five-fold cross-validation, and these cross-validation models were ensembled for hold-out testing and external validation on the Transcanadian Study and OCEAN Challenge datasets. Reporting followed the TRIPOD+AI checklist. The best-performing model used the H-optimus-0 foundation model, with five-class balanced accuracies of 89%, 97%, and 74% in the test sets. Normalisations and augmentations aided the performance of the ImageNet-pretrained ResNets, but these were still outperformed by 13 of the 14 foundation models. Hyperparameter tuning the downstream classifiers improved performance by a median 1.9% balanced accuracy, with many improvements being statistically significant. Histopathology foundation models offer a clear benefit to ovarian cancer subtyping, improving classification performance to a degree where clinical utility is tangible, albeit with an increased computational burden. Such models could provide a second opinion to histopathologists diagnosing challenging cases and may improve the accuracy, objectivity, and efficiency of pathological diagnoses overall. Code is made available at https://github.com/scjjb/Ovarian_Features.
keywords:
@@ Introduction
Histological subtyping is an essential component of the diagnostic process, but it can be challenging. From an individual slide, pathologists only exhibit concordance on an ovarian cancer diagnosis around 80% of the time [6]. In cases of uncertainty, a pathologist may request ancillary tests (such as P53 immunohistochemistry) or seek a second opinion from a gynaecological subspeciality expert, which incurs associated logistical and financial burdens. With increasing cancer rates [1] and growing complexity in diagnostic testing, histopathology services are increasingly struggling to meet demand globally. For example, most histopathology departments in the UK regularly resort to outsourcing work or hiring temporary staff [7], despite the UK being one of the countries with the most pathologists per capita [8]. Any delays resulting from demand outstripping diagnostic resources risk catastrophic impacts on patient outcomes, with a four-week delay in cancer treatment being associated with an approximately 10% increased mortality rate among patients [9].
Conceptually, artificial intelligence (AI) may offer clinical value by providing a second opinion to histopathologists, streamlining the diagnostic process and offering additional support when subspecialty experts are not readily available [10]. However, AI models for ovarian cancer diagnosis have yet to demonstrate clinical utility, with most research being small-scale prototyping [11] without regulatory approval for clinical use in Europe or the United States [12]. AI for ovarian cancer subtyping has constituted a small field of research where, aside from our work [13, 14, 15], research has almost exclusively been published by a single group [16, 17, 18, 19, 20, 21, 22, 23]. While the accuracy of such models has increased over time, the best models still only achieve around 80% accuracy [15, 21, 22, 23, 24, 25] and lack sufficient real-world testing.
@@
Recently, many researchers have attempted to create histopathology ‘foundation models’, using self-supervised learning (SSL) techniques to generate broad histopathological feature sets which are not specific to a single organ/cancer type. These approaches have grown rapidly, from tens of thousands of WSIs used to train models with tens of millions of parameters in 2022 and early 2023 [34, 35, 36, 37, 38, 39] to millions of WSIs [40, 41, 42] and billions of parameters more recently [25, 43]. Foundation models have typically been based on vision transformers (ViTs), utilizing the impressive scalability of transformers seen across many fields, most notably with large language models [44, 45]. Histopathology foundation models have exhibited impressive performance across diverse tasks [37, 46, 47, 48] including ovarian cancer subtyping [15, 24, 25], although analyses have been relatively shallow, without thorough hyperparameter tuning and rigorous statistical comparison of downstream models. Consequently, it is unclear whether models were applied optimally (especially those exhibiting sub-optimal performance), and whether the differences between them were significant. Furthermore, many analyses have been conducted using single-centre data, limiting the assessment of models’ generalisability.
@@
In this study, we present the most comprehensive validation conducted to date comparing feature extraction methods for ovarian cancer subtyping, including three ImageNet-pretrained feature extractors and fourteen histopathology foundation models. The analysis includes rigorous hyperparameter tuning and evaluations through five-fold cross-validation, hold-out testing, and external validations, and was conducted with the largest collection of ovarian cancer WSIs used in any AI validation study to date. We further investigate whether the classification performance of the ImageNet-pretrained ResNet50 features can match those of the foundation models through stain normalisation, tissue augmentation, or different tissue detection techniques.
Materials and Methods
Ovarian Carcinoma Histopathology Data
@@
Carcinoma Subtype Training WSIs (Patients) Hold-out WSIs (Patients) Transcanadian [33] WSIs (Patients) OCEAN [49] WSIs
High-Grade Serous (HGSC) 1266 (308) 20 (7) 30 (30) 217
Low-Grade Serous (LGSC) 92 (21) 20 (6) 9 (9) 42
Clear Cell (CCC) 198 (45) 20 (7) 20 (20) 94
Endometrioid (EC) 209 (38) 20 (5) 11 (11) 119
Mucinous (MC) 99 (22) 20 (5) 10 (10) 41
Total 1864 (434) 100 (30) 80 (80) 513